If a camera lens gives the proper exposure for a photograph at a shutter speed of 1/200 s at an f-stop of f/2.80, the proper shutter speed at f77.92 is 


  • 1/125 s. 

  • 1/100 s. 

  • 1/50.0 s

  • 1/25.0 s
Answer #1

So as we know Shutter speed is proportional to the frame rate so,

E\propto F

So,

\frac{E_1}{E_3}=\frac{F_1}{F_2}

So we get,

\frac{1/200s}{E_2}=\frac{f/2.8}{f/7.92}

So we get, minimum shutter speed for

E_2 = 1/50s